# Leveraging Discrete Function Decomposability for Scientific Design

### v1

@@ 4.2 Comparison on protein property predictors
Here we anchor our experiments on $f(x)$ fit to real protein property datasets. We use four datasets describing aggregation (Amyloid), virus viability (AAV), binding affinity (GB1), and cell toxicity (TDP-43) details in App. A.5). For each dataset, we used AlphaFold3-predicted structures (abramson24_af3) on the wild-type sequence to obtain a 3D structure, from which we constructed a contact graph by thresholding the distance between pairs of residues with threshold $t$. Following brookes22_sparsity; romero13_GP_protein; voigt02_recombination, we use a threshold of $t=4.5\text{\AA }$. We interpret this contact map as a graph adjacency matrix, from which we algorithmically construct a junction tree (lauritzen88_JT). This defines the decomposition graph needed for DADO, but we must also fit the corresponding component functions on the protein assay-labeled data. To do so, we employ an MLP-based predictive model that strictly enforces the decomposition defined by the junction tree. In particular, we use an MLP that takes as input the sequence, and, critically, also a bit-vector specifying active variables. One function evaluation requires calling this MLP for every node and edge function and summing. We used 5-fold cross validation to sweep through hidden layers of size ($[16,16]$, $[128,16]$, or $[128,128,16]$), learning rate ($0.001$ or $0.0001$), and number of training iterations ($5,000$ or $50,000$), to choose hyperparameters with the lowest cross-validation mean-squared error. Finally, we trained the model with those hyper-parameters using all data available. The result, a predictive model, $f(x)$, can be evaluated on any sequence $x\in\mathcal{X}$. We initialized the search distribution by training on the 1,000 sequences in the protein property dataset with the lowest assay labels, for 1,000 iterations, with a learning rate of $2\times 10^{-3}$.
Both methods draw $K=1000$ samples from their search distribution at each iteration. For three out of four datasets (Amyloid, AAV, and GB1), DADO reaches areas of high $f(x)$ much more quickly than the decomposition-unaware EDA (Fig. 3). In the fourth case, TDP-43, both methods converge rapidly and there is no noticeable difference in performance, despite the protein being longer than the others and hence the search space combinatorially larger. This can happen if $f$ has many local optima with near-optimal values, making it very easy to optimize. Our experiments on a few more proteins also support the conclusion that DADO either outperforms the naive EDA or is comparable (Fig. A3). When we decrease $K$ to $100$ (Fig. A4), we observe an improvement over the EDA for some saturated problems (both methods converged to roughly the same $f(x)$ at $K=1000$), including TDP-43. The choice of $K$ can substantially affect how well a DO algorithm optimizes $f$; we find that using a larger $K$ tends to benefit both methods and that DADO performs better than or comparably to the decomposition-unaware EDA for all $K$ tested (Sec. A.6.2). Beyond $K$, when a practitioner chooses a decomposition for their predictive model, there will likely be choices that are more or less decomposed with similar predictive accuracy. All other things being equal, we expect that a more decomposed model will be easier to optimize, especially for a decomposition-aware method. In a case study on GB1, we test different choices of the contact threshold, $t$, and observe that a more decomposed model (smaller $t$) can be optimized much more quickly and effectively by DADO for a relatively small decrease in accuracy. (Sec. A.6.3; Fig. A5).
@@ 5 Discussion
We have proposed a new method for distributional optimization that can leverage arbitrary decomposability of the function being optimized. We have shown that DADO works as expected on synthetic problems–namely, better than a naive, decomposition-unaware EDA. We have also demonstrated DADO’s potential for practical utility on the problem of protein design. Importantly, it is not necessary that a problem strictly adhere to the specified decomposition in order to be useful. For instance, we found that by adjusting the contact threshold applied to AlphaFold3-predicted 3D structures, we could obtain a range of decomposed functions, including some with increased decomposability but comparable predictive accuracy. This result adds to observations in the BO community suggesting that approximate or even random decompositions can be helpful for optimization (e.g., kandasamy15_additive-BO; ziomek23_random-decompositions_BO). Obtaining similarly useful decomposability in other domains will require further investigation; however, as the real world typically is structured, we expect many areas will be amenable to doing so. Estimating decomposability from labeled data is an active area of research (poelwijk16_wht; grudzien24_fgm; park24_rfa).
It will be insightful to conduct future work interrogating the use of DADO with generalizations of the naive EDA, such as conditioning an unconditional generative model with an inaccurate property predictor in order to perform AI-guided design while accounting for the fact that $f(x)$ is not truly known everywhere due to finite data (brookes19_cbas). This particular problem may require distilling an unconditional model into one that respects the decomposition, although alternative strategies could prove useful. Additionally, DADO, like naive EDAs, is in principle agnostic to whether the design space is discrete or continuous. Investigating decomposability for continuous scientific design problems, such as those in trabucco22_design-bench, is an interesting direction for future work. Exploring extensions of policy optimization from RL (sutton98_rl-book) in the context of DADO could also be fruitful, such as using learned value functions. Finally, use of DADO within BO for optimization of the acquisition function is an exciting direction.
@@ 6 Reproducibility statement
We specify our experimental setup, hyperparameters, model architectures, and algorithm details in Sec. 4. Datasets used and their sources are detailed in App. A.5. Code for our method and experiments is available publicly on GitHub: https://github.com/james-bowden/DADO.
@@ Acknowledgments
We thank Kuba Grudzien and members of the Listgarten group for helpful discussions regarding function decomposition, Hanlun Jiang and Peter Ma for discussions regarding scientific applications, and Toby Kreiman, Bear Xiong, and Aileen Zhang for feedback on the manuscript.
@@ A.1 Different ways of writing a function decomposition
However, $f(x)=C_{1}(\hat{x}_{1})+C_{2}(\hat{x}_{2}),\dots,C_{\kappa}(\hat{x}_{\kappa})$ as written has no interpretation in terms of a graph, which is necessary for message-passing and for factorizing the search distribution (Sec. 2.1). To bridge this gap, we introduce a graph in which nodes are design variables and edges specify which design variables directly interact to influence $f$. Users of DADO may also specify the decomposition at this level; we expect this to be helpful when domain knowledge is available in the negative form, i.e., when the user can begin with the fully-connected graph and remove edges between design variables they think don’t directly interact. This graph can then be automatically transformed into a junction tree [lauritzen88_JT], which is what our algorithm actually takes as input. That is, DADO takes an undirected junction tree, $\mathcal{T}:=(\mathcal{N},\mathcal{E})$, with nodes, $\mathcal{N}$, and edges, $\mathcal{E}$. Each node is a set of design variable indices, such that indexing $x$ with this set yields a new meta-variable, $\tilde{x}_{i}$, correspondingly exactly to a node in the junction tree, whereas the old meta-variables $\hat{x}_{i}$ had no such interpretation. $f(x)$ must be decomposed instead according to $\mathcal{T}$.
@@ A.1.1 Connection to epistasis
Function decomposition corresponds intimately to notions of the epistatic landscape for a protein property function (e.g.poelwijk16_whtwu82_pottsotwinowski13_truncated-epistatic-regressionlipsh24_epistasis-protein-design). An “epistatic expansion” of $f$,
@@
An expansion of this sort can be thought of as a spectral decomposition of $f$ into components of lower and higher frequencies, and can be computed for small landscapes using a discrete Fourier transform. A function with primarily lower-frequency components would be smooth and easy to optimize whereas a function with higher-frequency components would be rugged and difficult to find the global optimum of. One can imagine a worst-case scenario, that of $L$-order epistasis, of which a needle-in-a-haystack function is an example. In general, full epistatic landscapes are intractable to compute for design spaces of practical sizes, but estimation of a subset of the terms is an active area of research (e.g.otwinowski13_truncated-epistatic-regression, park24_rfa) due to beliefs that natural proteins tend to exhibit sparse and mostly lower-order epistasis.
@@ A.2 Decomposition-unaware EDA derivation
Given an objective, $f(x)$, to maximize, a decomposition-unaware EDA (or “naive” EDA) transforms the original problem into a distributional optimization problem as follows [brookes20_eda_em]:
@@
For the equivalence to hold, the search distribution, $p_{\theta}(x)$, must be capable of representing a point mass on $x^{*}$. Intuitively, one can think of an EDA as having a search distribution, $p_{\theta}(x)$, that acts as a spotlight on the design space, which is iteratively moved toward areas of the space with high $f(x)$ in expectation. Modern EDAs parameterize the search distribution with neural network generative models, such as a Variational Autoencoder (kingma13_vae; e.g.,  brookes19_cbas).
Procedurally, after having initialized the search distribution, $p_{\theta^{0}}(x)$, this naive EDA then iterates through these three steps either $N$ times or until some convergence criteria is met [larranaga01_eda, brookes20_eda_em]:
@@
where the entropy term can be dropped because it has no dependence on $\theta$ and division by $Z$ can be dropped without changing the objective’s maximizer. The sample approximation of the remaining terms is exactly step 3. Assuming an exact update (i.e., infinite samples such that the expectation is evaluated exactly, $p_{\theta}(x)$ has sufficient capacity to represent $p^{*}_{n}(x)$, and the KL divergence reaches 0), the resulting search distribution at iteration $n$ is $p^{*}_{n}(x)\propto f(x)^{n}p_{\theta^{0}}(x)$. As $n\rightarrow\infty$, this distribution concentrates all of its mass at the global maxima of $f$ and as a result, $f(x^{*})=\mathbb{E}_{p_{\theta}(x)}[f(x)]$. The derivation in brookes18_dbas is perhaps closest to this one, though their search distribution improvement operator is motivated by Bayes’ rule. It can be generalized to the derivation we give by simply writing $p^{*}_{n}(x)\propto C(x)\cdot p_{\theta^{n}}(x)$ for some $C$. Bayes’ rule yields a special case of this improvement operator, namely one where the multiplier is $C(x)=\frac{p(S\mid x)}{\sum_{x}p_{\theta^{n}}(x)p(S\mid x)}$. We choose not to interpret $f(x)$ as a normalized distribution in our derivation, though one can definitely do so.
@@
In practice, the weighted MLE problem in step 3 is typically solved with gradient descent; we will choose to use only a fixed number of gradient steps, which can be theoretically justified through the equivalence of EDAs to Expectation-Maximization [brookes20_eda_em].
A.3 DADO derivation
@@
A.3.1 Classical message-passing for non-distributional optimization
Although classical message-passing has been largely used for probabilistic inference on probabilistic graphical models [pearl88_bp_textbook, shah_14_mit_inference-algorithms_message-passing], it can also be used for exact optimization of a function, $x^{*}=\arg\max\nolimits_{x}f(x)$ [vlassis04_maxplus]. In particular, message-passing can be used to find a global maximum of a function defined on an undirected junction tree, $\mathcal{T}$, by making use of its topology for optimal time-complexity. In particular, one takes the junction tree, which is undirected, and roots it to obtain a directed tree, which induces a hierarchy among the meta-variables from root to leaves. Each node is responsible for accumulating information from all nodes in its sub-tree and then passing this information on to its parent. Consequently, the root node receives information from the entire tree, which is sufficient to set its variables in a globally optimal manner. Then, starting with the root, each parent communicates its variables’ optimal settings to its children, which can in turn set their variables optimally, and so forth.
@@ A.3.1 Classical message-passing for non-distributional optimization
A.3.2 From classical message passing to distributional optimization
@@
Each factor is weighted by its corresponding value function, enabling it to coordinate with all its descendant factors despite their being updated separately. The whole DO is tied together at the top by the root factor. Our resulting algorithm (Alg. 1) fits into the three EDA steps: (1) designs are sampled from $p_{\theta^{n}}(x)$, (2) weights, here each $Q$-function instead of $f(x)$, are computed, and (3), each search distribution factor receives its own independent weighted maximum likelihood update based on these weights (steps 1 and 2 illustrated in Fig. 1b). These are repeated until convergence, or for some fixed number of iterations. The factor updates are coupled only through the $Q$-functions, the messages across edges. $\{Q^{\theta^{n}}_{i}\}_{i\in\mathcal{N}}$ are only valid while $\theta$ is close to $\theta^{n}$, meaning one must balance how many gradient steps are taken before drawing new samples. If too many gradient steps are taken without resampling, a factor may be changing its parameters to collaborate with another factor which has already changed its behavior. It’s common for EDAs to include an additional hyperparameter $W(\cdot)$, a monotonic shaping function applied to the weights, to alter optimization dynamics [brookes20_eda_em], so we include it in Alg.1( 1 on lines 16 and 18. Choosing $W$ to be the identity recovers our derivation above. Notice that in place of the classical message-passing traceback equations, DADO simply performs sequential conditional sampling from its search distribution. Interestingly, a very similar algorithm can be derived without using a lower bound on the value functions (Eq. 32) if one adds an entropy-maximizing term to the initial DO objective (Sec. A.4).
A.4 Alternate DADO Derivation: Maximum Entropy Objective
If, by comparison to Sec. A.3, we instead seek to solve a closely related objective which additionally maximizes the search distribution’s entropy, we can arrive at a similar method by leveraging the probabilistic graphical modeling (PGM) framework in levine18_control-as-inference. This view will provide another intuition for the relationship between classical value functions and distributional value functions approximated with Monte-Carlo samples. It also highlights the core components of our conceptual framework, which are shared even by derivations starting from different objectives. In other words, one need not use our specific method as outlined in Alg. 1 to reap the benefits of decomposition-aware distributional optimization, and variations may have desirable properties and/or simply perform better in certain scientific design settings. This derivation will assume that you’ve already read key parts of Sec. 2A.3 and are familiar with the function decomposition, search distribution factorization, and value functions.

### v2

@@ 4 Experimental Results
Here we anchor our experiments on $f(x)$ fit to real protein datasets. We use four protein property datasets, AAV, Amyloid, Gcn4, and TDP-43 (details in App. A.6). For each dataset, we used AlphaFold3-predicted structures (Abramson et al., 2024) on the wild-type sequence to obtain a 3D structure, from which we constructed a contact graph by thresholding the distance between pairs of residues with threshold $t$. Following Brookes et al. (2022); Romero et al. (2013); Voigt et al. (2002), we use a threshold of $t=4.5\text{\AA }$. We interpret this contact map as a graph adjacency matrix, from which we algorithmically construct a junction tree (Lauritzen and Spiegelhalter, 1988). This defines the topology needed for DADO (Fig. A4), and then we fit the component functions on the protein assay-labeled data (details in Sec. A.8.3). We initialized the search distribution by training on the 1,000 sequences with lowest $f(x)$, for 1,000 iterations, with a learning rate of $2\times 10^{-3}$.
All methods drew $K=1000$ samples from their search distribution at each iteration. For all four proteins, DADO converges to designs with higher $f(x)$ than the baselines both visually and by a statistical test (Fig. 3). For three of the four (Amyloid, Gcn4, and TDP-43), the seed-averaged 95% confidence intervals of DADO and the best baseline at the final iteration do not overlap. We plotted $\log(c-f(x))$ on the y-axis to make this difference clear visually; plots with $f(x)$ on the y-axis are included in the appendix (Fig. A8). Finally, in Sec. A.9.5, we observe that our choice of AlphaFold3-based topology is robust to mild modifications as far as holdout performance is concerned.
@@ 5 Discussion
We have proposed a new method for distributional optimization that can leverage arbitrary decomposability of the function being optimized. We have shown that it works as expected on synthetic problems–namely, better than a naive EDA that is not aware of the decomposition. We have also demonstrated the potential for practical utility on the problem protein design. Importantly, it is not necessary that a problem strictly adhere to the specified decomposition in order to be useful. Specifically, we showed that using the heuristic of thresholding AlphaFold3 estimated contacts, we can obtain a range of decomposed functions, including some that maintain predictive accuracy while providing a level of decomposability helpful for DADO. Obtaining similarly useful decomposability in other domains will require further investigation; however, as the real world typically is structured, we expect many areas will be amenable to doing so. Estimating decomposability from labeled data is an active area of research (Poelwijk et al. (2016); Grudzien et al. (2024); Park et al. (2024)).
It will be insightful to conduct future work interrogating the use of DADO with generalizations of the standard EDA, such as conditioning an unconditional generative model with an inaccurate property predictor for AI-guided design, which should enable us to design while accounting for the fact that $f(x)$ is not truly known everywhere in the space due to finite data. This particular problem may require distilling an unconditional model into one that respects the decomposition, although alternative strategies could prove useful. Finally, use of DADO within Bayesian Optimization for optimization of the acquisition function is an exciting direction.
@@ 6 Reproducibility Statement
We specify exhaustively our experimental setup, hyperparameters, model architectures, and algorithm details in Sec. 4. Datasets used and their sources are detailed in Sec. A.6. Code for our method and experiments is available publicly on GitHub: https://github.com/james-bowden/DADO.
@@ Acknowledgments
We thank Kuba Grudzien for helpful discussions regarding function decomposition, Hanlun Jiang and Peter Ma for discussions regarding scientific applications, Toby Kreiman and Bear Xiong for feedback on the manuscript, and Aileen Zhang for feedback on the figures.
@@ A.1 Different Ways of Writing a Function Decomposition
However, $f(x)=C_{1}(\hat{x}_{1})+C_{2}(\hat{x}_{2}),\dots,C_{\kappa}(\hat{x}_{\kappa})$ as written has no interpretation in terms of a graph, which is necessary for message-passing and for factorizing the search distribution (Sec. 2.1). To bridge this gap, we introduce a graph in which nodes are design variables and edges specify which design variables directly interact to influence $f$. Users of DADO may also specify the decomposition at this level; we expect this to be helpful when domain knowledge is available in the negative form, i.e., when the user can begin with the fully-connected graph and remove edges between design variables they think don’t directly interact. This graph can then be automatically transformed into a junction tree (Lauritzen and Spiegelhalter, 1988), which is what our algorithm actually takes as input. That is, DADO takes an undirected junction tree, $\mathcal{T}:=(\mathcal{N},\mathcal{E})$, with nodes, $\mathcal{N}$, and edges, $\mathcal{E}$. Each node is a set of design variable indices, such that indexing $x$ with this set yields a new meta-variable, $\tilde{x}_{i}$, correspondingly exactly to a node in the junction tree, whereas the old meta-variables $\hat{x}_{i}$ had no such interpretation. $f(x)$ must be decomposed instead according to $\mathcal{T}$.
@@ A.1.1 Connection to epistasis
Function decomposition corresponds intimately to notions of the epistatic landscape for a protein property function (e.g.Poelwijk et al. (2016)Wu (1982)Otwinowski and Nemenman (2013)Lipsh-Sokolik and Fleishman (2024)). An “epistatic expansion” of $f$,
@@
An expansion of this sort can be thought of as a spectral decomposition of $f$ into components of lower and higher frequencies, and can be computed for small landscapes using a discrete Fourier transform. A function with primarily lower-frequency components would be smooth and easy to optimize whereas a function with higher-frequency components would be rugged and difficult to find the global optimum of. One can imagine a worst-case scenario, that of $L$-order epistasis, of which a needle-in-a-haystack function is an example. In general, full epistatic landscapes are intractable to compute for design spaces of practical sizes, but estimation of a subset of the terms is an active area of research (e.g.Otwinowski and Nemenman (2013); Park et al. (2024)) due to beliefs that natural proteins tend to exhibit sparse and mostly lower-order epistasis.
@@ A.2 Decomposition-Unaware EDA Derivation
Given an objective, $f(x)$, to maximize, a decomposition-unaware EDA (or “naive” EDA) transforms the original problem into a distributional optimization problem as follows (Brookes et al., 2020):
@@
For the equivalence to hold, the search distribution, $p_{\theta}(x)$, must be capable of representing a point mass on $x^{*}$. Intuitively, one can think of an EDA as having a search distribution, $p_{\theta}(x)$, that acts as a spotlight on the design space, which is iteratively moved toward areas of the space with high $f(x)$ in expectation. Modern EDAs parameterize the search distribution with neural network generative models, such as a Variational Autoencoder (Kingma and Welling (2013); e.g.,  Brookes et al. (2019)).
Procedurally, after having initialized the search distribution, $p_{\theta^{0}}(x)$, this naive EDA then iterates through these three steps either $N$ times or until some convergence criteria is met (Larrañaga and Lozano, 2001; Brookes et al., 2020):
@@
where the entropy term can be dropped because it has no dependence on $\theta$ and division by $Z$ can be dropped without changing the objective’s maximizer. The sample approximation of the remaining terms is exactly step 3. Assuming an exact update (i.e., infinite samples such that the expectation is evaluated exactly, $p_{\theta}(x)$ has sufficient capacity to represent $p^{*}_{n}(x)$, and the KL divergence reaches 0), the resulting search distribution at iteration $n$ is $p^{*}_{n}(x)\propto f(x)^{n}p_{\theta^{0}}(x)$. As $n\rightarrow\infty$, this distribution concentrates all of its mass at the global maxima of $f$ and as a result, $f(x^{*})=\mathbb{E}_{p_{\theta}(x)}[f(x)]$. The derivation in Brookes and Listgarten (2018) is perhaps closest to this one, though their search distribution improvement operator is motivated by Bayes’ rule. It can be generalized to the derivation we give by simply writing $p^{*}_{n}(x)\propto C(x)\cdot p_{\theta^{n}}(x)$ for some $C$. Bayes’ rule yields a special case of this improvement operator, namely one where the multiplier is $C(x)=\frac{p(S\mid x)}{\sum_{x}p_{\theta^{n}}(x)p(S\mid x)}$. We choose not to interpret $f(x)$ as a normalized distribution in our derivation, though one can definitely do so.
@@
In practice, the weighted MLE problem in step 3 is typically solved with gradient descent; we will choose to use only a fixed number of gradient steps, which can be theoretically justified through the equivalence of EDAs to Expectation-Maximization (Brookes et al., 2020).
A.3 Algorithmic Sketch: From EDA to DADO
In this section, we describe how infusing a naive EDA (Sec. A.2) with awareness of a known decomposition of $f(x)$ stands to make the EDA more statistically efficient.
Intuitively, one can think of an EDA as having a search distribution, $p_{\theta}(x)$, that acts as a spotlight on the design space, which is iteratively moved toward areas of the space with high $f(x)$ in expectation. In a decomposition-unaware EDA, after having initialized the search distribution, $p_{\theta^{0}}(x)$, with initial parameters, $\theta^{0}$, the EDA proceeds by iterating through these three steps for $n=1\dots N$ or until some convergence criteria has been met (Sec. A.2):
  1. 1.
  2. 2.
  3. 3.
Now that a standard EDA is fresh in our minds, we can consider what it might mean to infuse it with knowledge of how $f(x)$ is decomposed, and how this might prove useful. Before discussing the general case, here we first build intuition by considering a simpler example, one that is “fully”-decomposable rather than only soft-decomposable. That is, consider $f(x)=C_{1}(\hat{x}_{1})+C_{2}(\hat{x}_{2})$ where the two meta-variables $\hat{x}_{1}$ and $\hat{x}_{2}$ do not share any original variables, $x_{i}$, corresponding to a junction tree without any edges, only independent nodes. In this setting, we can break the EDA down into two independent EDAs, each with reduced dimensionality compared to the original EDA. Specifically, we can employ a fully-factorized search distribution (“fully” again denoting that no variable appears in more than one factor), $p_{\theta}(x)=p_{\theta_{1}}(\hat{x}_{1})p_{\theta_{2}}(\hat{x}_{2})$. We modify the three steps of the EDA as follows:
  1. 1.
  2. 2.
  3. 3.
Earlier we described decomposability as $f(x)=C_{1}(\hat{x}_{1})+C_{2}(\hat{x}_{2}),\dots,C_{\kappa}(\hat{x}_{\kappa})$ to help provide some intuition. However, to develop DADO it will be more convenient to equivalently describe the decomposability with a junction tree defined by nodes that contain multiple design variables and undirected edges that connect the nodes. We define this junction tree decomposition in Sec. 2.1 and explain its connection to the above decomposability in Sec. A.1. It suffices to say here that design variables $x_{i}$ that occurred uniquely in one meta-variable $\hat{x}_{j}$ now sit in a single node, $\tilde{x}_{k}$; variables that previously occurred in multiple meta-variables now appear in multiple adjacent nodes. With the idea of a junction tree in hand, we are now set to complete our intuitive overview of DADO, which will use a factorized search distribution matching the topology of the junction tree. Namely, we will convert the undirected junction tree into a directed, acyclic graph (DAG) by rooting it and then directing the edges outward from the root. This operation will yield a search distribution that can be written as $p_{\theta}(x)=p_{\theta}(\tilde{x}_{r})\sum_{i}{p_{\theta}(\tilde{x}_{i}\mid\tilde{x}_{\text{p}(i)})}$, for root node $\tilde{x}_{r}$ and non-root nodes, $\tilde{x}_{i}$, with parents, $\tilde{x}_{\text{p}(i)}$. Note that each non-root node has exactly one parent in a DAG. Each search distribution factor is a conditional probability distribution that will be parameterized by a neural network such as an autoregressive model.
  1. 1.
  2. 2.
  3. 3.
For a formal derivation of DADO, see Sec. 2.1 or Sec. A.4.
A.4 DADO Derivation
@@
A.4.1 Classical message-passing for non-distributional optimization
Although classical message-passing has been largely used for probabilistic inference on probabilistic graphical models (Pearl, 1988; Shah, 2014), it can also be used for exact optimization of a function, $x^{*}=\arg\max\nolimits_{x}f(x)$ (Vlassis et al., 2004). In particular, message-passing can be used to find a global maximum of a function defined on an undirected junction tree, $\mathcal{T}$, by making use of its topology for optimal time-complexity. In particular, one takes the junction tree, which is undirected, and roots it to obtain a directed tree, which induces a hierarchy among the meta-variables from root to leaves. Each node is responsible for accumulating information from all nodes in its sub-tree and then passing this information on to its parent. Consequently, the root node receives information from the entire tree, which is sufficient to set its variables in a globally optimal manner. Then, starting with the root, each parent communicates its variables’ optimal settings to its children, which can in turn set their variables optimally, and so forth.
@@ A.4.1 Classical message-passing for non-distributional optimization
A.4.2 From classical message passing to distributional optimization
@@
Each factor is weighted by its corresponding value function, enabling it to coordinate with all its descendant factors despite their being updated separately. The whole DO is tied together at the top by the root factor. Our resulting algorithm (Alg. 1) fits into the three EDA steps: (1) designs are sampled from $p_{\theta^{n}}(x)$, (2) weights, here each $Q$-function instead of $f(x)$, are computed, and (3), each search distribution factor receives its own independent weighted maximum likelihood update based on these weights (steps 1 and 2 illustrated in Fig. 1b). These are repeated until convergence, or for some fixed number of iterations. The factor updates are coupled only through the $Q$-functions, the messages across edges. $\{Q^{\theta^{n}}_{i}\}_{i\in\mathcal{N}}$ are only valid while $\theta$ is close to $\theta^{n}$, meaning one must balance how many gradient steps are taken before drawing new samples. If too many gradient steps are taken without resampling, a factor may be changing its parameters to collaborate with another factor which has already changed its behavior. It’s common for EDAs to include an additional hyperparameter $W(\cdot)$, a monotonic shaping function applied to the weights, to alter optimization dynamics (Brookes et al., 2020), so we include it in Alg.1( 1 on lines 16 and 18. Choosing $W$ to be the identity recovers our derivation above. Notice that in place of the classical message-passing traceback equations, DADO simply performs sequential conditional sampling from its search distribution. Interestingly, a very similar algorithm can be derived without using a lower bound on the value functions (Eq. 23) if one adds an entropy-maximizing term to the initial DO objective (Sec. A.5).
A.5 Alternative DADO Derivation: Maximum Entropy Objective
If, by comparison to Sec. A.4, we instead seek to solve a closely related objective which additionally maximizes the search distribution’s entropy, we can arrive at a similar method by leveraging the probabilistic graphical modeling (PGM) framework in Levine (2018). This view will provide another intuition for the relationship between classical value functions and distributional value functions approximated with Monte-Carlo samples. It also highlights the core components of our conceptual framework, which are shared even by derivations starting from different objectives. In other words, one need not use our specific method as outlined in Alg. 1 to reap the benefits of decomposition-aware distributional optimization, and variations may have desirable properties and/or simply perform better in certain scientific design settings. This derivation will assume that you’ve already read key parts of Sec. 2/Sec. A.4 and are familiar with the function decomposition, search distribution factorization, and value functions.